OpenStudy (anonymous):

What is the range of the graph of y = 5(x – 2)2 + 7? Answer y less than or equal to 7 y greater than or equal to 7 y less than or equal to 2 y greater than or equal to 2

OpenStudy (anonymous):

y = 5(x – 2)^2 + 7?

OpenStudy (saifoo.khan):

sorry idk this one. =(

OpenStudy (anonymous):

its ok hunni :)

OpenStudy (anonymous):

The second

OpenStudy (anonymous):

5(x – 2)2? means 5*(x-2)*2?

OpenStudy (saifoo.khan):

it's squared.. thta's obv.

OpenStudy (saifoo.khan):

that's*

OpenStudy (anonymous):

Ah alrighty.

OpenStudy (anonymous):

y = 5(x – 2)2 + 7 the min x-2 can = is 0 at x=2 y=5(2-2)^2+7 y=5(0)^2+7 y=7 so it's the second one.